Please help me solve this equation: −9 + 8v = 7 + 4v
THE STEP HERE IS MISSING
4v = 16
v = 4
Please fill in the step and solve the problem! thank you so much :)

Answer by math-vortex(648)

Hi, there--
.
−9 + 8v = 7 + 4v
THE STEP HERE IS MISSING
4v = 16
v = 4
.
The Missing Step:
.
Actually, two things happen in the missing step. Overall, we are combining like terms. We want all the variable terms on the left hand side and all the constant terms on the right hand side.
.
Start with
-9 + 8v = 7 + 4v
.
Add 9 from both sides and subtract 4v from both sides.
-9 + 8v + 9 - 4v = 7 + 4v + 9 - 4v (This is your missing step)
.
Notice that -9 + 9 = 0 so you are left with 8v - 4v = 4v on the left hand side.
On the right hand side, 4v - 4v = 0, so you are left with 7 + 9 = 16 on the right hand side.
This takes you to the next step :
4v = 16
.
Don't be too concerned that you couldn't figure out what happened in the missing step. There is a LOT going on.
.
Feel free to email if you have questions about this explanation.
.
Ms.Figgy
math.in.the.vortex@gmail.com